A welded joint, as shown in Fig. is subjected to an eccentric load of 2500 N.



Find the size of the weld, if the maximum shear stress in the weld is not to exceed



50 N/mm 2

1
Expert's answer
2022-04-06T01:02:03-0400

"A = \frac{P}{\sigma }"

"A = \frac{2500}{50}=50 mm^2"

"A = t \times l"

Throat thickness "t =\frac{50}{10}=5mm"

Size of weld "s = \frac{t}{k}"

"s = \frac{5}{0.7} =7.14 mm \approx 8 mm"